OpenStudy (anonymous):

Write the equation of the line in slope-intercept form with a slope of -6 and a y-intercept of 9.

OpenStudy (anonymous):

we are given that slope m= -6 and y intercept i.e. c= 9 the eq of a line in slope - intercept form is given as: y = mx + c i.e. y = -6x + 9 (by substituting regarding values) i.e. 6x + y = 9 and this us the required eq of line in the given form.

OpenStudy (anonymous):

Slope intercept form is y = mx + b so the equation you need is : y = -6x + 9